OpenStudy (anonymous):

HW Help: Using the following equation, find the center and radius: x^2 - 4x + y^2 + 8y = -4 Select one: a. The center is located at (-2, -4), and the radius is 4. b. The center is located at (2, -4), and the radius is 4. c. The center is located at (-2, -4), and the radius is 16. d. The center is located at (2, -4), and the radius is 16.

OpenStudy (anonymous):

b. The center is located at (2, -4), and the radius is 4.

OpenStudy (anonymous):

You know why?

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

\[ x^2 - 4x + y^2 + 8y = -4 \\ (x^2 - 4x +4) -4 +( y^2 + 8y+16) -16 = -4 \\ (x-2)^2 + (y+4)^2=-4+16+4=16=4^2\\ (x-2)^2 + (y+4)^2==4^2\\ \]

OpenStudy (anonymous):

That's it all worked out right?

OpenStudy (anonymous):

Yes